The most common human blood group is O which is estimated to be 45% in US population. If you randomly sample 5 people from the US population and test their blood group,
(a) what is the probability that exactly 3 of them have type O blood?
(b) what is the mean and standard deviation of the number of people with blood type O?

Answer by Boreal(15235)

mean is np=5*0.45=2.25
variance is np(1-p)=2.25*0.55=1.2375
sd is sqrt(V)=1.11
exactly 3 is 5C3*0.45^3*0.55^2=0.2757